OpenStudy (anonymous):

when looking at this: 2 / (x+h-3)(x-3) why is A=2, B=1, and C=-3 ?

OpenStudy (asnaseer):

the question does not make sense - where did A, B and C come from?

OpenStudy (anonymous):

those are the answers

OpenStudy (anonymous):

i wanna know how you get them

OpenStudy (asnaseer):

there must be more information in the question that you have missed out

OpenStudy (anonymous):

i made the question

OpenStudy (anonymous):

by looking at the first thing how do you get the A, B, and C?

OpenStudy (asnaseer):

sorry - I can't help here - it really makes no sense to me.

OpenStudy (anonymous):

this was supposed to be a simpler question -_-

OpenStudy (anonymous):

how would you find A, B, and C by looking at the first thing? (phrased better?)

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

If f(x) = x2 - 6x + 1, then f(1) = -4, f(2) = -7, f(3) = -8, .... p(x) = (x - h)q(x) + p(h); they give us the expression ... Find numbers a, b, c, and d such that f(x) ... and q integers having no common integral divisor greater that 1, then p must be an integral .

OpenStudy (anonymous):

"2 is to 3 as 10 is to 15." Because 2 is two thirds of 3, just as 10 is two thirds of 15. ... c) x : y = 7 : 10. x is seven tenths of y. .... h) In a class of 40 pupils, 25 got a B.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

For example: In a class of 40 pupils, 25 got a B.

OpenStudy (anonymous):

In order to learn calculus, you have got to be able to do algebra. .... In the previous example, where h = 3, k = 5, and m = 2, this leads to a very quick solution: ... Suppose you had y = 2x - 1 as one equation and y = 0.4x + 3 as the other. ... Of course, a, b, and c need not be nice neat symbols, as I have here, but each might be .

OpenStudy (anonymous):

start the power at 1 and increase by 1 each time, getting up to n. ... But how are we going to solve for A, B, and D. We cannot plug in numbers to ... If we did, we would see that A = 2, B = 2, C = 3, D = -4, E = -3, F = 0, G = 3, H = -3, and I = -2.

OpenStudy (anonymous):

ur losing me... my brain is not as advanced as urs lol

OpenStudy (anonymous):

k? then wat grade r u in

OpenStudy (anonymous):

so i can simple it up

OpenStudy (anonymous):

im in college but im terrible at math

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

?